OpenStudy (anonymous):

What is the surface area of a circle that has a diameter of 14 inches?

OpenStudy (anonymous):

You mean area?

OpenStudy (anonymous):

49 pi

OpenStudy (anonymous):

Or 154 Square inches

OpenStudy (anonymous):

A = (pi) r² A = (3) (49) Area is around 150 inches squared for a more accurate answer we can use 22/7 for pi. A =22/7 times 49 / 1 you can cancel the 7 in the denominator off with a factor of 7 from the 49 in the numerator: A = 22 / 1 times 7 / 1 A =22 times 7 A =154

OpenStudy (anonymous):

these are the answer chorces A. 49.86 in2 B. 153.86 in2 C. 196.86 in2 D. 616.86 in2

OpenStudy (anonymous):

153.86

OpenStudy (anonymous):

154 s. in. formula would be 

OpenStudy (anonymous):

so the answer is b

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ya it is